192 J of heat was supplied to a thermodynamic system and its internal energy was

found to decrease by 22 J. Calculate the amount of associated work and state

whether the work was done by the system or done on the system?


1
Expert's answer
2021-01-15T06:43:11-0500

according to First law of thermodynamics du = dq+dw ; du = change in internal energy, dq= heat change, dw= work done

as heat supplied to the system so dq=+192 J

as internal energy chnge is negetive deu to decrease in internal energy , du = -22 J

dw = du - dq =-22 - (+192 ) = -22-192 = -214 J ;

as dw is negetive according to IUPAC convention work done by the system. here system done work with the heat absorbed and also utilise its internal energy so total work done by the system become -214 J.